A geometry textbook gives the equation of a parabola as y = x 2, where x
and y are measured in inches. How can this equation be dimensionally correct?
1
Expert's answer
2021-08-30T01:54:32-0400

To make the equation "y=x^2" dimensionally correct, there must be a constant of either side with a dimension that balances both sides. Let the constant be k.

Therefore,

"y=kx^2\\
Dimensionally,\\
L=kL^2\\
k=L^{1-2}=L^{-1}"